OpenStudy (anonymous):

solve x^2+x+3=0 (mod 101)

OpenStudy (anonymous):

Use the quadratic formula man! You keep posting the same type of questions over and over. The formula is: (-b +/- SQRT(b^2-4ac)) / 2a

OpenStudy (anonymous):

This is a congruence equation. (number theory)

OpenStudy (anonymous):

hmm what do you mean mod 101? :-)

OpenStudy (anonymous):

modulo

OpenStudy (anonymous):

hmm |101| ?

OpenStudy (anonymous):

basicly it is just for what x is x^2+x+3 divisible by 101

OpenStudy (anonymous):

(there will be 0 or infinite many solutions)

OpenStudy (anonymous):

oh okay cool stuff

OpenStudy (anonymous):

But i dont know how to solve this one... First idea was to complete the square but that leaves the integers and that sucks :-)

OpenStudy (anonymous):

hmm x^2 + x + 3 --------------- = k right? 101 k for integers now (x + 1/2)^2 + 3 - 1/4 = k * 101 (x + 1/2)^2 = k*101 - 11/4 hmm maybe like this

OpenStudy (anonymous):

\[x + \frac{1}{2} = \sqrt{k*101 - 11/4}\]

OpenStudy (anonymous):

now for x to be valid \[k*101 > 11/4\]

OpenStudy (across):

\[x=\pm\sqrt{101n-\frac{11}{4}}-\frac{1}{2},\]\[\forall n\in\mathbb{Z}.\]What about this?

OpenStudy (anonymous):

\[k > \frac{11}{4 \times 101}\]

OpenStudy (anonymous):

x can only be integer as well!

OpenStudy (anonymous):

oh I didn't knew about that sorry it's my first time solving these

OpenStudy (anonymous):

number theory is all about integers

OpenStudy (anonymous):

and divisibility also only makes sense with the integers

OpenStudy (anonymous):

hmm k > 11/101*4 hmm and this is a small value so k must be valid for nearly every positive integer now

OpenStudy (anonymous):

what I did may not be right but worth a thought: multiply by 34 to get 34(x^2+x)+102=0 (mod 101) which is the same as 34(x^2+x)+1=0 (mod 101) Or we can multiply by 67 (67*3=201) which is -1 mod 101 67(x^2+x)-1=0 (mod101)

OpenStudy (anonymous):

Ok I will ask my seminar leader on Friday :-)